OpenStudy (anonymous):

what is a number between 0.005 and 0.0005

OpenStudy (bahrom7893):

there are infinitely many numbers in that interval

OpenStudy (bahrom7893):

0.0005111111111111111 is in there, so is 0.0005111111 or 0.00051111112,etc

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

If you are looking at the middle number between the two. The ans should be 0.00275. 0.005-0.0005=0.0045 0.0045/2=0.00225 0.00225+0.0005=0.00275

OpenStudy (mertsj):